OpenStudy (anonymous):

When simplifying the expression (2/x )+(2/x-1)+(2/x+2). The denominator is x(x-1)(x+2), what is the numerator? please give a solid answer and explnation! Oi vay I hate math!

OpenStudy (anonymous):

2*(x-1)(x+2) + 2*x*(x+2) + 2*x*(x-1)

OpenStudy (precal):

\[2(x^2-1x-2)+2x^2+4x+2x^2-2x\] \[2x^2-2x-4+2x^2+4x+2x^2-2x\] \[6x^2-4\]

OpenStudy (precal):

using Stiwan's solution and simplifying it further

OpenStudy (anonymous):

The reason for this are the elementary rules of fractional arithmetic. If you need further explanations I suggest you visit http://en.wikipedia.org/wiki/Fraction_%28mathematics%29#Addition The rule used here is\[\frac{ a }{ b } + \frac{ c }{ d } = \frac{ ad + cb }{ bd }\]